xloverofmine:

factor the trinomial. 6x^2 + x - 40

jhonyy9:

do you can calcule the roots of this quadratic ?

jhonyy9:

P(x)= ax^2 +bx +c and x_1 , x_2 the roots of P(x) so in this case you can write P(x) = a(x - x_1)(x - x_2) the factorized form

Verifiedkylaagang:

What grade is this?/

surjithayer:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @surjithayer product of 6 and -40=-240 16-15=1 \[16\times15=-240\] \[6x^2+x-40\] \[=6x^2+(16-15)x-40\] \[=6x^2+16x-15x-40\] =2x(3x+8)-5(3x+8) =(3x+8)(2x-5) \(\color{#0cbb34}{\text{End of Quote}}\)